OpenStudy (anonymous):

simplify (-7)^5 divided by (-7)^3

jimthompson5910 (jim_thompson5910):

When you divide exponential expressions with the same base, you subtract the exponents

jimthompson5910 (jim_thompson5910):

The result of the subtraction will be the final exponent

jimthompson5910 (jim_thompson5910):

Ex: 2^9 divided by 2^5 = 2^(9-5) = 2^4

OpenStudy (anonymous):

49

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

i would of been better if people just gave an answer instead of explaining and wasting time

OpenStudy (anonymous):

seriously ! -.- lol thanks why i thank you for saving me time :D

OpenStudy (anonymous):

thats *

OpenStudy (anonymous):

welcome